OpenStudy (anonymous):

During a clothing store’s Bargain Days, the regular price for T-shirts is discounted by $5. There is a state sales tax of 5%, and the $5 discount is applied before the sales tax is calculated. a. Write an expression that shows the regular price r of a T-shirt minus the $5 discount. b. Write a rule for the function p(r) that expresses the final price p of a T-shirt with the discount applied and sales tax added. c. How much would you pay during Bargain Days for a shirt regularly priced at $15.50?

OpenStudy (anonymous):

Let p= the regular price of a t-shirt in dollars. Let s= the sales price of a t-shirt in dollars. Let r= the sales tax rate as decimal. So you have Part A: \[s=p-5.00\] Part B; \[p(r)=(p-5.00)(r+1)\] Part C: \[p(r)=(15.50-5.00)(1+.05)\] Which gives us $15.75

OpenStudy (anonymous):

thanks(: